# Symmetrical Flow Matching: Unified Image Generation, Segmentation, and Classification with Score-Based Generative Models

### v2

@@ 1 Introduction
Comprehending semantic content is a key challenge in computer vision. Classification (krizhevsky2012imagenet; he2016deep; dosovitskiy2020image) and segmentation (ronneberger2015u; xie2021segformer; cheng2022masked) allow models to analyze and structure images, while generative modeling enables the synthesis of new content (radford2015unsupervised; ho2020denoising; rombach2022high). Ideally, a unified framework would bridge these tasks, allowing models to both interpret and generate images in a two-way manner. It is conceivable that the ability to accurately comprehend and disentangle visual structures facilitates the generation of more semantically coherent and visually realistic images. Conversely, strong generative capabilities may aid in learning more expressive representations of images, as generating plausible content requires an implicit understanding of object relationships, textures, and context (he2022masked). These forward and backward relationships suggest that advances in one direction could naturally benefit the other, which fuels the pursuit of models that integrate both understanding and synthesis within a single cohesive framework.
Most existing vision-only approaches treat these tasks individually. For classification, the extracted features are used in fully connected layers to obtain probability predictions for each class, while for segmentation, the features are used to train decoders for dense predictions. On the other hand, generative models, such as generative adversarial networks (GANs) (karras2021alias), diffusion models (karras2022elucidating), score matching models (song2020score), and Flow Matching models (lipman2022flow), synthesize images from a prior distribution. Although recent work has explored diffusion models for classification (li2023your) and segmentation (wu2024medsegdiffv2), these adaptations introduce significant limitations: classification is slow due to the need for iterative sampling across all possible classes, and segmentation frameworks are restricted to generating masks, lacking the ability to map back to realistic images.
Recent works, such as SemFlow (wang2024semflow) and DepthFM (gui2024depthfm), have sought to unify image generation with semantic segmentation and depth estimation within a single generative framework. However, these models still suffer from the following key limitations: (1) they do not perform classification, restricting their applicability; (2) the image quality remains inferior to that of purely generative models; and (3) the models enforce a strict one-to-one mapping between segmentation or depth masks and images, requiring them to have the same number of channels, which limits their flexibility.
@@ 2.1 Flow Matching
Flow Matching (FM) (lipman2022flow; liu2022flow) is a generative modeling framework that learns a velocity field to transform a source distribution into a target distribution through a continuous flow. By parameterizing this field with a neural network, the model becomes a Neural ODE (chen2018neural), allowing efficient sampling via numerical integration. It has advanced state-of-the-art performance in diverse applications, including image (esser2024scaling) and audio generation (vyas2023audiobox), as well as protein modeling (huguet2024sequence) and robotics (black2024pi_0). FM generalizes Continuous Normalizing Flows (CNFs) (chen2018neural; grathwohl2018ffjord) by eliminating the need for simulation during training, thereby making it computationally efficient. It also provides a unifying perspective on generative modeling, encompassing diffusion models, which can be seen as a special case where the probability path is defined via stochastic differential equations (lipman2024flow).
@@ 2.2 Generative Classifiers
Several seminal works (hinton2007recognize; ranzato2011deep) have emphasized the importance of modeling the data distribution to enhance discriminative feature learning. Early approaches trained deep belief networks (hinton2006fast) to encode image data as latent representations, which were then used for recognition tasks. More recent advances in generative modeling have demonstrated the ability to learn efficient representations for global prediction tasks (he2022masked). In addition, generative models have been shown to improve adversarial robustness and calibration (huang2020neural). However, most prior work either jointly trains generative and discriminative models, or fine-tunes generative representations for downstream tasks. Diffusion Classifier (li2023your) specifically investigates the effectiveness of using diffusion models as image classifiers, albeit with severe inference constraints.
@@ 2.3 Semantic Segmentation
Semantic segmentation aims to assign a semantic label to each pixel in an image. Conventional approaches rely on discriminative models, combining a strong feature extraction backbone with a task-specific decoder head for mask prediction (cheng2021per; cheng2022masked; ding2023mevis). Recent work has explored diffusion models for segmentation (amit2021segdiff; wang2024explore; gu2024diffusioninst), typically leveraging them as feature extractors within a discriminative framework. A key motivation behind SemFlow is the interpretation that diffusion models struggle to align their stochastic nature with the deterministic requirements of semantic segmentation. For this reason, SemFlow introduces rectified flows. In contrast, SymmFlow embraces the probabilistic nature of segmentation, accounting for inter-observer variability.
@@ 2.4 Semantic Image Generation
Semantic image generation, the inverse of semantic segmentation, focuses on generating realistic images from semantic layouts (li2019diverse; liu2019learning; lv2022semantic). Existing approaches generally fall into two categories. (1) GAN-based models (zhu2017unpaired; isola2017image; wang2018high), although many of these methods struggle with mode collapse and produce only unimodal outputs. (2) Diffusion models, which treat semantic generation as a conditional generation task where semantic masks act as control signals (wang2022semantic). Some methods further integrate additional conditioning to enhance coherence (zhang2023adding). However, these approaches often adopt asymmetric architectures with unidirectional generators, thereby complicating the unification of semantic segmentation and image generation.
@@ 4 Experiments
$S$ DML-CSR (zheng2022decoupled) 1 77.8
SegFace (narayan2025segface) 1 81.6
DeeplabV2 (caesar2018coco) 1 33.2
MaskFormer (cheng2021per) 1 37.1
SegFormer (xie2021segformer) 1 46.7
SIS pix2pixHD (wang2018high) 1 54.7 / 0.529 111.5 / —
SPADE (park2019semantic) 1 42.2 / 0.487 33.9 / —
SC-GAN (wang2021image) 1 19.2 / 0.395 18.1 / —
BBDM (li2023bbdm) 200 21.4 / 0.370
ControlNet (zhang2023adding) 20 24.0 / 0.528 36.6 / 0.671
SDM (wang2022semantic) 1000 18.8 / 0.422 15.9 / 0.518
SCDM (ko2024stochastic) 250 17.4 / 0.418 15.3 / 0.519
SCP-Diff (gao2024scp) 800 11.3 / —
Both SemFlow (wang2024semflow) 25 *69.4 *35.7 32.6 / 0.393 *90.0 / 0.685
@@ 4.1 Datasets
Semantic Segmentation and Generation: The model is evaluated using COCO-Stuff (caesar2018coco) and CelebAMask-HQ (lee2020maskgan), which contain 171 and 19 classes, respectively. Images and semantic masks are resized and cropped into $\times512$ pixels. A detailed analysis of the datasets is provided in Appendix A. Classification: We evaluate our model on MNIST (deng2012mnist) and CIFAR-10 (krizhevsky2010cifar) for classification, leveraging these low-resolution datasets to assess fundamental capabilities.
@@ 4.5 Baseline Comparison
Semantic Segmentation and Generation: For semantic segmentation on CelebAMask-HQ, DML-CSR (zheng2022decoupled) and SegFace (narayan2025segface) are adopted as baselines. On COCO-Stuff, DeeplabV2 (caesar2018coco), MaskFormer (cheng2021per), and SegFormer (xie2021segformer) serve as representative methods. For semantic image synthesis, the evaluation includes pix2pixHD (wang2018high), SPADE (park2019semantic), SC-GAN (wang2021image), BBDM (li2023bbdm), ControlNet (zhang2023adding), SDM (wang2022semantic), SCDM (ko2024stochastic), and SCP-Diff (gao2024scp). SemFlow (wang2024semflow), a prior model aiming to unify both segmentation and synthesis, is also evaluated. Since official checkpoints and semantic segmentation evaluation scripts for SemFlow were not publicly available, the model was retrained and results were recomputed. The scripts are included in SymmFlow’s repository for reproducibility. Classification: For the classification task on CIFAR-10, the proposed model is compared to the Diffusion Classifier (li2023your). Additionally, the quality of the generated images is assessed by benchmarking against a standard FM setup.
@@ 4.6 Implementation Details
For pixel-level implementations (e.g., classification), the U-Net architecture introduced in Guided Diffusion (dhariwal2021diffusion) is employed. For latent-space models, the pre-trained VAE from Stable Diffusion (rombach2022high) is used for image encoding and decoding, alongside the U-Net backbone from Stable Diffusion 2.1. To ensure compatibility with SymmFlow, the number of input channels in the first layer and output channels in the final layer of the U-Net are doubled. Further details on hyperparameters and computational resources are provided in Appendix B.
@@ 6 Limitations & Future Work
One limitation of the current approach is the overall model size. Although SymmFlow uses far fewer inference steps than typical diffusion models, it relies on a large pre-trained Stable Diffusion U-Net backbone. Reducing the computational burden by distilling the model into a one-step variant is a natural next step. Moreover, a one-step formulation could facilitate the integration of more expressive conditioning mechanisms, enabling effective text-based control and the use of more powerful architectures such as MMDiT (esser2024scaling). Additionally, fine-tuning the VAE decoder to better align with semantic masks could improve segmentation accuracy on fine-grained details, such as small or occluded regions.
Future work includes extending classification evaluation from the current proof-of-concept stage to datasets such as Food-101 (bossard2014food), ImageNet-1K (deng2009imagenet), and ObjectNet (barbu2019objectnet), while also refining the semantic label encoding strategy to improve conditioning. Furthermore, evaluating the model performance on depth estimation tasks, similar to DepthFM, would further demonstrate its versatility. Beyond segmentation, the bi-directionality of the model presents opportunities for applications such as image editing.
@@ Appendix A Datasets
This section provides details on the datasets used for classification and semantic segmentation and generation. For classification, we use MNIST (deng2012mnist) and CIFAR-10 (krizhevsky2010cifar), while for segmentation, we evaluate on CelebAMask-HQ (lee2020maskgan) and COCO-Stuff (caesar2018coco). The datasets are summarized in Table 6.
Table 6: Summary of the benchmark datasets.
@@ Appendix B Implementation Details
For pixel-space implementations on MNIST and CIFAR-10, we use the U-Net architecture proposed by Dhariwal et al. (dhariwal2021diffusion), which has been widely adopted for diffusion-based generative modeling. The architecture consists of a series of residual blocks, self-attention layers, and group normalization, enabling effective denoising and feature extraction across multiple resolutions. Table 7 presents the hyperparameters used for training these two models. To ensure reproducibility, we will also make the code publicly available.
Table 7: Hyperparameters used for training each model.
@@
Table 8: Hyperparameters used for training each model.
@@ Appendix C Image Generation Latency
Model Latency (ms)
ControlNet (zhang2023adding) 5,483.8
SemFlow (wang2024semflow) 2,140.1
SymmFlow (Proposed) 2,156.9

### v3

@@ 1 Introduction
Comprehending semantic content is a key challenge in computer vision. Classification (Krizhevsky, Sutskever, and Hinton 2012; He et al. 2016; Dosovitskiy et al. 2020) and segmentation (Ronneberger, Fischer, and Brox 2015; Xie et al. 2021; Cheng et al. 2022) allow models to analyze and structure images, while generative modeling enables the synthesis of new content (Radford, Metz, and Chintala 2015; Ho, Jain, and Abbeel 2020; Rombach et al. 2022). Ideally, a unified framework would bridge these tasks, allowing models to both interpret and generate images in a two-way manner. It is conceivable that the ability to accurately comprehend and disentangle visual structures facilitates the generation of more semantically coherent and visually realistic images. Conversely, strong generative capabilities may aid in learning more expressive representations of images, as generating plausible content requires an implicit understanding of object relationships, textures, and context (He et al. 2022). These forward and backward relationships suggest that advances in one direction could naturally benefit the other, which fuels the pursuit of models that integrate both understanding and synthesis within a single cohesive framework.
Most existing vision-only approaches treat these tasks individually. For classification, the extracted features are used in fully connected layers to obtain probability predictions for each class, while for segmentation, the features are used to train decoders for dense predictions. On the other hand, generative models, such as generative adversarial networks (GANs) (Karras et al. 2021), diffusion models (Karras et al. 2022), score matching models (Song et al. 2020), and Flow Matching models (Lipman et al. 2022), synthesize images from a prior distribution. Although recent work has explored diffusion models for classification (Li et al. 2023a) and segmentation (Wu et al. 2024), these adaptations introduce significant limitations: classification is slow due to the need for iterative sampling across all possible classes, and segmentation frameworks are restricted to generating masks, lacking the ability to map back to realistic images.
Recent works, such as SemFlow (Wang et al. 2024b) and DepthFM (Gui et al. 2024), have sought to unify image generation with semantic segmentation and depth estimation within a single generative framework. However, these models still suffer from the following key limitations: (1) they do not perform classification, restricting their applicability; (2) the image quality remains inferior to that of purely generative models; and (3) the models enforce a strict one-to-one mapping between segmentation or depth masks and images, requiring them to have the same number of channels, which limits their flexibility.
@@ 2.1 Flow Matching
Flow Matching (FM) (Lipman et al. 2022; Liu, Gong, and Liu 2022) is a generative modeling framework that learns a velocity field to transform a source distribution into a target distribution through a continuous flow. By parameterizing this field with a neural network, the model becomes a Neural ODE (Chen et al. 2018), allowing efficient sampling via numerical integration. It has advanced state-of-the-art performance in diverse applications, including image (Esser et al. 2024) and audio generation (Vyas et al. 2023), as well as protein modeling (Huguet et al. 2024) and robotics (Black et al. 2024). FM generalizes Continuous Normalizing Flows (CNFs) (Chen et al. 2018; Grathwohl et al. 2018) by eliminating the need for simulation during training, thereby making it computationally efficient. It also provides a unifying perspective on generative modeling, encompassing diffusion models, which can be seen as a special case where the probability path is defined via stochastic differential equations (Lipman et al. 2024).
@@ 2.2 Generative Classifiers
Several seminal works (Hinton 2007; Ranzato et al. 2011) have emphasized the importance of modeling the data distribution to enhance discriminative feature learning. Early approaches trained deep belief networks (Hinton, Osindero, and Teh 2006) to encode image data as latent representations, which were then used for recognition tasks. More recent advances in generative modeling have demonstrated the ability to learn efficient representations for global prediction tasks (He et al. 2022). In addition, generative models have been shown to improve adversarial robustness and calibration (Huang et al. 2020). However, most prior work either jointly trains generative and discriminative models, or fine-tunes generative representations for downstream tasks. Diffusion Classifier (Li et al. 2023a) specifically investigates the effectiveness of using diffusion models as image classifiers, albeit with severe inference constraints.
@@ 2.3 Semantic Segmentation
Semantic segmentation aims to assign a semantic label to each pixel in an image. Conventional approaches rely on discriminative models, combining a strong feature extraction backbone with a task-specific decoder head for mask prediction (Cheng, Schwing, and Kirillov 2021; Cheng et al. 2022; Ding et al. 2023). Recent work has explored diffusion models for segmentation (Amit et al. 2021; Wang et al. 2024a; Gu, Chen, and Xu 2024), typically leveraging them as feature extractors within a discriminative framework. A key motivation behind SemFlow is the interpretation that diffusion models struggle to align their stochastic nature with the deterministic requirements of semantic segmentation. For this reason, SemFlow introduces rectified flows. In contrast, SymmFlow embraces the probabilistic nature of segmentation, accounting for inter-observer variability.
@@ 2.4 Semantic Image Generation
Semantic image generation, the inverse of semantic segmentation, focuses on generating realistic images from semantic layouts (Li, Zhang, and Malik 2019; Liu et al. 2019; Lv et al. 2022). Existing approaches generally fall into two categories. (1) GAN-based models (Zhu et al. 2017; Isola et al. 2017; Wang et al. 2018), although many of these methods struggle with mode collapse and produce only unimodal outputs. (2) Diffusion models, which treat semantic generation as a conditional generation task where semantic masks act as control signals (Wang et al. 2022). Some methods further integrate additional conditioning to enhance coherence (Zhang, Rao, and Agrawala 2023). However, these approaches often adopt asymmetric architectures with unidirectional generators, thereby complicating the unification of semantic segmentation and image generation.
@@ 4 Experiments
$S$ DML-CSR (Zheng et al. 2022) 1 77.8
SegFace (Narayan, Vs, and Patel 2025) 1 81.6
DeeplabV2 (Caesar, Uijlings, and Ferrari 2018) 1 33.2
MaskFormer (Cheng, Schwing, and Kirillov 2021) 1 37.1
SegFormer (Xie et al. 2021) 1 46.7
SIS pix2pixHD (Wang et al. 2018) 1 54.7 / 0.529 111.5 / —
SPADE (Park et al. 2019) 1 42.2 / 0.487 33.9 / —
SC-GAN (Wang et al. 2021) 1 19.2 / 0.395 18.1 / —
BBDM (Li et al. 2023b) 200 21.4 / 0.370
ControlNet (Zhang, Rao, and Agrawala 2023) 20 24.0 / 0.528 36.6 / 0.671
SDM (Wang et al. 2022) 1000 18.8 / 0.422 15.9 / 0.518
SCDM (Ko et al. 2024) 250 17.4 / 0.418 15.3 / 0.519
SCP-Diff (Gao et al. 2024) 800 11.3 / —
Both SemFlow (Wang et al. 2024b) 25 *69.4 *35.7 32.6 / 0.393 *90.0 / 0.685
@@ 4.1 Datasets
Semantic Segmentation and Generation: The model is evaluated using COCO-Stuff (Caesar, Uijlings, and Ferrari 2018) and CelebAMask-HQ (Lee et al. 2020), which contain 171 and 19 classes, respectively. Images and semantic masks are resized and cropped into $\times512$ pixels. A detailed analysis of the datasets is provided in Appendix A. Classification: We evaluate our model on MNIST (Deng 2012) and CIFAR-10 (Krizhevsky, Nair, and Hinton 2010) for classification, leveraging these low-resolution datasets to assess fundamental capabilities.
@@ 4.5 Baseline Comparison
Semantic Segmentation and Generation: For semantic segmentation on CelebAMask-HQ, DML-CSR (Zheng et al. 2022) and SegFace (Narayan, Vs, and Patel 2025) are adopted as baselines. On COCO-Stuff, DeeplabV2 (Caesar, Uijlings, and Ferrari 2018), MaskFormer (Cheng, Schwing, and Kirillov 2021), and SegFormer (Xie et al. 2021) serve as representative methods. For semantic image synthesis, the evaluation includes pix2pixHD (Wang et al. 2018), SPADE (Park et al. 2019), SC-GAN (Wang et al. 2021), BBDM (Li et al. 2023b), ControlNet (Zhang, Rao, and Agrawala 2023), SDM (Wang et al. 2022), SCDM (Ko et al. 2024), and SCP-Diff (Gao et al. 2024). SemFlow (Wang et al. 2024b), a prior model aiming to unify both segmentation and synthesis, is also evaluated. Since official checkpoints and semantic segmentation evaluation scripts for SemFlow were not publicly available, the model was retrained and results were recomputed. The scripts are included in SymmFlow’s repository for reproducibility. Classification: For the classification task on CIFAR-10, the proposed model is compared to the Diffusion Classifier (Li et al. 2023a). Additionally, the quality of the generated images is assessed by benchmarking against a standard FM setup.
@@ 4.6 Implementation Details
For pixel-level implementations (e.g., classification), the U-Net architecture introduced in Guided Diffusion (Dhariwal and Nichol 2021) is employed. For latent-space models, the pre-trained VAE from Stable Diffusion (Rombach et al. 2022) is used for image encoding and decoding, alongside the U-Net backbone from Stable Diffusion 2.1. To ensure compatibility with SymmFlow, the number of input channels in the first layer and output channels in the final layer of the U-Net are doubled. Further details on hyperparameters and computational resources are provided in Appendix B.
@@ 6 Limitations & Future Work
One limitation of the current approach is the overall model size. Although SymmFlow uses far fewer inference steps than typical diffusion models, it relies on a large pre-trained Stable Diffusion U-Net backbone. Reducing the computational burden by distilling the model into a one-step variant is a natural next step. Moreover, a one-step formulation could facilitate the integration of more expressive conditioning mechanisms, enabling effective text-based control and the use of more powerful architectures such as MMDiT (Esser et al. 2024). Additionally, fine-tuning the VAE decoder to better align with semantic masks could improve segmentation accuracy on fine-grained details, such as small or occluded regions.
Future work includes extending classification evaluation from the current proof-of-concept stage to datasets such as Food-101 (Bossard, Guillaumin, and Van Gool 2014), ImageNet-1K (Deng et al. 2009), and ObjectNet (Barbu et al. 2019), while also refining the semantic label encoding strategy to improve conditioning. Furthermore, evaluating the model performance on depth estimation tasks, similar to DepthFM, would further demonstrate its versatility. Beyond segmentation, the bi-directionality of the model presents opportunities for applications such as image editing.
@@ Appendix A Datasets
This section provides details on the datasets used for classification and semantic segmentation and generation. For classification, we use MNIST (Deng 2012) and CIFAR-10 (Krizhevsky, Nair, and Hinton 2010), while for segmentation, we evaluate on CelebAMask-HQ (Lee et al. 2020) and COCO-Stuff (Caesar, Uijlings, and Ferrari 2018). The datasets are summarized in Table 6.
@@ Appendix B Implementation Details
For pixel-space implementations on MNIST and CIFAR-10, we use the U-Net architecture proposed by Dhariwal et al. (Dhariwal and Nichol 2021), which has been widely adopted for diffusion-based generative modeling. The architecture consists of a series of residual blocks, self-attention layers, and group normalization, enabling effective denoising and feature extraction across multiple resolutions. Table 7 presents the hyperparameters used for training these two models. To ensure reproducibility, we will also make the code publicly available.
@@
Table 7: Hyperparameters used for training each model.
@@
Table 8: Hyperparameters used for training each model.
@@ Appendix C Image Generation Latency
Model Latency (ms)
ControlNet (Zhang, Rao, and Agrawala 2023) 5,483.8
SemFlow (Wang et al. 2024b) 2,140.1
SymmFlow (Proposed) 2,156.9